# Privacy-Preserving Resilient Vector Consensus

### v1

@@ Abstract
This paper studies privacy-preserving resilient vector consensus in multi-agent systems against faulty agents, where normal agents can achieve consensus within the convex hull of their initial states while protecting state vectors from being disclosed. Specifically, we consider a modification of an existing algorithm known as Approximate Distributed Robust Convergence Using Centerpoints (ADRC), i.e., Privacy-Preserving ADRC (PP-ADRC). Under PP-ADRC, each normal agent introduces multivariate Gaussian noise to its state during each iteration. We first provide sufficient conditions to ensure that all normal agents’ states can achieve mean square convergence under PP-ADRC. Then, we analyze convergence accuracy from two perspectives, i.e., the Mahalanobis distance of the final value from its expectation and the Hausdorff distance based alteration of the convex hull caused by noise when only partial dimensions are added with noise. Then, we employ concentrated geo-privacy to characterize privacy preservation and conduct a thorough comparison with differential privacy. Finally, numerical simulations demonstrate the theoretical results.
@@ 1 Introduction
Enabling multi-agent systems to collaborate effectively in solving complex problems or accomplishing tasks has garnered significant interest in fields such as aviation, robotics, and others [1, 2, 3]. In many application scenarios, such as distributed machine learning, multi-robot systems, and platoon, vector consensus is frequently used, aiming to make all agents’ state vectors reach agreement under a predefined rule. However, in multi-agent systems, there may be faulty or adversarial agents [^1] In the subsequent text, both are collectively referred to as faulty agents. against the rule to break the consensus of the system or drive the consensus to an unsafe value. To maintain the safe consensus of normal agents, the concept of resilient vector consensus has been developed. Resilient vector consensus means that all normal agents can achieve vector consensus where the final value lies within the convex hull of the initial states of normal agents despite the presence of faulty agents. Under resilient vector consensus, many distributed systems can work correctly under complex and changing circumstances. Therefore, resilient vector consensus is a very important concept in distributed fields.
In most consensus/resilient algorithms, agents directly send their states to their neighbors. However, if the transmitted states are intercepted by a malicious entity, this could lead to the disclosure of agents’ valuable information, resulting in a privacy leakage. For example, in distributed machine learning applications like the financial and medical sectors, the exchange of grid information during the training process may lead to the leakage of sensitive data from training sets. The sensitive data may include bank customers’ account balances, transaction records, as well as patients’ diagnosis results and medication histories. Such disclosure can potentially result in financial fraud and serious breaches of patient privacy, indicating that privacy preservation is a crucial task in the resilient vector consensus.
@@
Due to the scarcity of communication and computation resources in distributed systems and the high demand of encryption algorithms on these resources, differential privacy has become the primary choice for privacy protection in consensus algorithms. Typically, in the resilient vector consensus, we aim to protect the initial state vectors of the normal agents without the existence of a trusted control center, where local differential privacy (local-DP) [4] is suitable to be utilized. Different from centralized differential privacy (central-DP), local-DP protects information during transmission from being intercepted, unrelated to differential attacks, and thus does not involve the concept of neighboring inputs. Local-DP also ensures that any pair of inputs generate similar outputs. However, the initial state range for each agent can be very large, and requiring pairs of distant initial states to produce similar outputs would reduce the system’s utility [^2] Utility refers to the extent to which a system can still perform its original functions after the application of privacy-preserving mechanisms.. This motivates us to use the concept of geo-privacy, characterizing the Euclidean distance between agents as one of the metrics for privacy protection. Specifically, we propose to use the concept of concentrated geo-privacy [5], which is a generalization of differential privacy, to describe the privacy preservation of resilient vector consensus. Apart from a more precise description of privacy preservation, it also offers advantages such as Gaussian mechanism and advanced composition.
@@
Much attention has been paid to resilient consensus algorithms, which can be roughly classified into two categories. One is to detect and preclude faulty agents[6, 7, 8], while the other one is to try to find a safe value despite the presence of faulty ones[9, 10, 11, 12]. In this paper, we focus on the second one as it has more analytical results and related works regarding resilient scalar/vector consensus and privacy-preserving resilient scalar consensus are provided below.
@@
Resilient scalar consensus is designed for the normal agents to converge to a common state that lies between the maximum and minimum values of normal agents’ initial states. There are two frequently used ways: One way is to use “median”. In [13], Zhang et al. propose a median consensus algorithm, where each normal agent updates by using its own value and the median of the states from its neighborhood. As for the other way, most of the algorithms rely on the strategy of simply disregarding suspicious values. For example, in a series of algorithms named Mean-Subsequence-Reduced (MSR) algorithm [9, 10], the main idea is to discard a fixed number of largest and smallest values in each normal agent’s neighborhood. Conversely, in the modification of Weighted MSR (W-MSR) algorithm [11, 12], each normal agent only discards values that are larger or smaller than its own to maintain its state.
@@
Resilient vector consensus is ensured by that each normal agent must seek a point which is always located in the convex hull of its normal neighboring state vectors at each iteration. The existing work can be classified into three categories. The first category utilizes Tverberg partitions to compute Tverberg points. Specifically, it includes Byzantine vector consensus algorithm [14] and Approximate Distributed Robust Convergence algorithm [2]. The second category calculates the intersection of multiple convex hulls. This involves Algorithm $1$ proposed in [15] and the multidimensional approximate agreement algorithm described in [16]. The algorithm of the last category utilizes the concept of “centerpoint”, exemplified by “Approximate Distributed Robust Convergence Using Centerpoint” (ADRC) algorithm [17]. However, these algorithms are rather time-consuming, and most of them require approximation algorithms, which consequently reduce fault tolerance. Notably, among these algorithms, the ADRC [17] stands out with its better tolerance for the mature approximation algorithm. In each iteration, every normal agent computes the centerpoint (an extension of the median in higher dimensions) of its neighborhood and adjusts its position accordingly. Resilient vector consensus can be achieved as long as the communication topology and the number of faulty agents within the neighborhood meet certain requirements.
@@
Privacy-preserving resilient scalar consensus can be achieved by two general approaches, i.e., cryptography-based methods [18, 19] and noise-adding-based methods [18, 20]. Although cryptography-based algorithms have higher convergence accuracy compared to noise-adding-based ones, encryption, information exchange, and decryption are time-consuming processes and thus significantly decrease efficiency. Besides, the noise-adding-based ones are typically more flexible, as it allows for the adjustment of noise intensity and distribution based on specific application scenarios and privacy requirements. In [20], Fiore et al. proposed a Differentially Private MSR (DP-MSR) algorithm, adding decaying, zero-mean Laplace noise to the scalar states of normal agents. They also analyzed resilient scalar consensus in the sense of probability and differential privacy.
@@
Note that privacy-preserving resilient vector consensus is still an open issue. One simple way is to conduct a privacy-preserving resilient scalar consensus algorithm $d$ times, where $d$ is the dimension of states. However, this may cause the final value to fall outside the convex hull of the initial states of the normal agents [17]. Additionally, analyzing the performance of privacy-preserving resilient vector consensus by noise adding involves several key challenges: (i) The noise added makes the convergence non-deterministic and random, and the existing definition of resilient vector consensus is inapplicable. (ii) Due to the noise and resilience, we cannot derive the analytical solution of the final value, making it difficult to describe the accuracy of convergence. Additionally, high-dimensional convex hulls are inherently difficult to characterize and require specific values to obtain. After adding noise, the vertices and boundaries of the convex hull change randomly, making it challenging to characterize the convex hull after multiple iterations. (iii) At each iteration, each agent transmits and updates state information and the specific distribution of local state at each iteration is non-deterministic. We need to analyze the extent to which this information leaks the initial state, which is also a very difficult task.
@@
To solve the above challenges, we consider a modification of ADRC, named Privacy-Preserving ADRC (PP-ADRC), and provide rigorous theoretical analysis for convergence and privacy performance. Compared to our conference version [21], we use Gaussian noise instead of Laplace noise and conduct a more rigorous performance analysis, Meanwhile, we add the final value distribution analysis by characterizing the deviation of the final value from the expectation. Moreover, we use CGP instead of DP to characterize the privacy-preservation and compare it with DP comprehensively. We also add more simulations under a $3-$dimensional case to illustrate our theoretical results. The contributions of this paper are given as follows.
@@
  • We consider a modification of ADRC, named Privacy-Preserving ADRC (PP-ADRC), where each agent adds Gaussian noise to the local state vector for local interaction. We show sufficient conditions to ensure a resilient consensus of expectation.
@@
  • We employ $\rho$-concentrated geo-privacy to characterize privacy preservation without detailed distributions of the outputs. Through a detailed comparison with differential privacy, we demonstrate that $\rho-$concentrated geo-privacy can provide advantages for resilient vector consensus.
The organization of this work is as follows. We introduce preliminaries and problem formulation in Sections 2 and 3, respectively. Then, we analyze the convergence condition and final value in Section 4. The privacy analysis is presented in Section 5. Simulation results are presented in Section 6. Finally, conclusions and avenues for future research are outlined in Section 7.
@@ 2 PRELIMINARIES
Basic Notations: Throughout this paper, we denote by $\mathbb{R}^{d}$ the $d$-dimensional Euclidean space, by $\mathbb{R}^{m\times d}$ the space of all $m\times d$-dimensional matrices, and by $(\mathbb{R}^{m\times d})^{\mathbb{N}}$ the space of matrix valued sequences in $\mathbb{R}^{m\times d}$, where $\mathbb{N}$ denotes the set of natural numbers. The natural logarithm is denoted by $\log$. Let $\mathbf{A}$ be a matrix. Then, we denote by $[\mathbf{A}]_{i},[\mathbf{A}]^{j},[\mathbf{A}]_{ij},\mathbf{A}^{\top}$ the $i$th row, the $j$th column, the $(i,j)$th element, and the transpose of matrix $\mathbf{A}$. We then consider two matrices $\mathbf{A}$ and $\mathbf{B}$, and if $[\mathbf{A}]_{ij}\leq[\mathbf{B}]_{ij}$ holds for any pair of $i,j$, we say $\mathbf{A}\leq\mathbf{B}$. For a vector $x\in\mathbb{R}^{d}$, we denote by $x_{\operatorname{max}}$ and $x_{\operatorname{min}}$ maximum element and minimum element of it, respectively. Then, we consider a matrix consisting of $n$ row vectors $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]^{\top}$, where $x_{i}$ denotes the $i$th vector and $x_{i,k}$ stands for the $k$th element of the vector $x_{i}$. We denote by $\mathbf{1}$ a column vector of ones, and by $I_{d}$ the $d-$dimensional identity matrix. A matrix is row-stochastic if all its elements are non-negative and each row sums to $1$. For a given point set $C\subseteq\mathbb{R}^{d}$, we denote by $\text{conv}(C)$ the convex hull of the point set $C$, by $|C|$ the cardinality of $C$. We consider two functions $f(x)$ and $g(x)$, and $f(x)=\Omega(g(x))$ means that $f(x)\geq g(x)$. For two different vectors $x,x^{\prime}\in\mathbb{R}^{d}$, we represent by $\mathrm{dist}(x,x^{\prime})=||x-x^{\prime}||_{2}$ the Euclidean distance between $x$ and $x^{\prime}$. For two $n$-tuples of vectors $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]^{\top}$ and $\mathbf{x}^{\prime}=[x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{n}^{\prime}]^{\top}$, $\mathrm{dist}(\mathbf{x},\mathbf{x}^{\prime})$ stands for the maximum distance among all the vectors, which is $\max\limits_{i}\operatorname{dist}(x_{i},x_{i}^{\prime})$. In probability theory, given a random variable $z\in\mathbb{R}$, we denote by $\mathbb{E}(z)$ and $\operatorname{var}(z)$ the expectation and the variance of $z$, respectively. As for a random vector $z=[z_{1},\ldots,z_{d}]^{\top}\in\mathbb{R}^{d}$, the expectation of $z$ is $\mathbb{E}(z)=[\mathbb{E}(z_{1}),\ldots,\mathbb{E}(z_{d})]^{\top}$. For a topological space $\mathcal{T}$, we denote by $\mathcal{B}(\mathcal{T})$ the set of Borel subsets of $\mathcal{T}$, and $\mathrm{P}$ is the corresponding probability measure. For a zero mean $d-$dimensional Gaussian distribution, the probability density function is given by $G(x;\Sigma)=\frac{1}{\sqrt{(2\pi)^{d}|\Sigma|}}\mathrm{exp}\left(-\frac{1}{2}x^{\top}\Sigma^{-1}x\right)$.
@@
Privacy Notions: We first give two formal definitions of standard local differential privacy and then introduce CGP.
@@
($\varepsilon-$Differential Privacy): Given spaces $U$, $V$, and $\varepsilon\in\mathbb{R}\geq 0$, a randomized function $M:U\rightarrow V$ is $\varepsilon-$differentially private, if for any pair of inputs $x$, $x^{\prime}\in U$ and any $S\subseteq V$, we have $\mathrm{P}\{M(x)\in S\}\leq e^{\varepsilon}\mathrm{P}\{M(x^{\prime})\in S\}.$
@@
($(\varepsilon,\delta)-$Differential Privacy): Given spaces $U$, $V$, and $\varepsilon$, $\delta\in\mathbb{R}\geq 0$, a randomized function $M$ is $(\varepsilon,\delta)-$differentially private, if for any pair of inputs $x$, $x^{\prime}\in U$ and any $S\subseteq V$, $\mathrm{P}\{M(x)\in S\}\leq e^{\varepsilon}\mathrm{P}\{M(x^{\prime})\in S\}+\delta$ holds.
Note that in Definitions 3-4, the inputs $x$ and $x^{\prime}$ can be in the real domain that includes one or more scalars or vectors. The parameter $\varepsilon$ in both definitions characterizes the privacy protection strength, with smaller values of $\varepsilon$ providing stronger privacy protection. The unique difference between the two definitions is the relaxation term $\delta$, i.e., $(\varepsilon,\delta)$-DP means that $\varepsilon$-DP is achieved with a probability of at least $1-\delta$. It is noteworthy that the Gaussian mechanism can only achieve $(\varepsilon,\delta)$-DP.
Definitions 3-4 are quite strict for resilient vector consensus. As the range of possible initial states of an agent is typically large, ensuring two far apart initial states produce similar outputs would cause significant system utility reduction. To solve this issue, geo-privacy has been widely studied, which uses Euclidean distance as one metric for privacy protection. Supporting the Gaussian mechanism and better composition, CGP is selected here, which is established on Rényi divergence given below [22].
@@
Note that $D_{\alpha}(F\|G)$ is determined by the expectation of the $\alpha-1$ power of the ratio of two distributions at each point, quantifying the difference between two distributions. Meanwhile, larger $\alpha$ emphasize parts where the ratio is greater, while smaller values of $\alpha$ focus more on the overall average differences. The roles of logarithm and $\frac{1}{\alpha-1}$ are both intended to prevent numerical overflow in the results. Rényi divergence is monotonically non-decreasing with $\alpha$.
@@
(Concentrated Geo-Privacy): Given spaces $U$, $V$, and $\rho\in\mathbb{R}\geq 0$, a randomized function $M:U\rightarrow V$ is said to satisfy $\rho-$concentrated geo-privacy, if for any inputs $x,x^{\prime}\in U$ and all $\alpha>1$, it holds that
@@
In CGP, $\rho$ provides the upper bound on $D_{\alpha}(M(x)\|M(x^{\prime}))/\alpha/\mathrm{dist}(x,x^{\prime})^{2}$, characterizing the difference between the output distributions of $M$ for any pair of inputs, thereby quantifying the strength of privacy protection. A smaller $\rho$ indicates that $M$ is less sensitive to the inputs, thereby providing stronger privacy protection. Besides, the privacy should degrade gracefully as $\mathrm{dist}(x,x^{\prime})$ increases.
3 Problem Formulation
In this section, we first provide the network model of the multi-agent system with faulty agents and then propose the PP-ADRC algorithm. Finally, interesting problems are stated.

### v2

@@ Abstract
This paper studies privacy-preserving resilient vector consensus in multi-agent systems against faulty agents, where normal agents can achieve consensus within the convex hull of their initial states while protecting state vectors from being disclosed. Specifically, we consider a modification of an existing algorithm known as Approximate Distributed Robust Convergence Using Centerpoints (ADRC), i.e., Privacy-Preserving ADRC (PP-ADRC). Under PP-ADRC, each normal agent introduces multivariate Gaussian noise to its state during each iteration. We first provide sufficient conditions to ensure that all normal agents’ states can achieve mean square convergence under PP-ADRC. Then, we analyze convergence accuracy from two perspectives, i.e., the Mahalanobis distance of the final value from its expectation and the Hausdorff distance-based alteration of the convex hull caused by noise when only partial dimensions are added with noise. Then, we employ concentrated geo-privacy to characterize privacy preservation and conduct a thorough comparison with differential privacy. Finally, numerical simulations demonstrate the theoretical results.
@@ 1 Introduction
Enabling multi-agent systems to collaborate effectively in solving complex problems or accomplishing tasks has garnered significant interest in fields such as aviation, robotics, and others [1, 2, 3]. In many application scenarios, such as distributed machine learning, multi-robot systems, and platoon, vector consensus is frequently used, aiming to make all agents’ state vectors reach agreement under a predefined rule. However, in multi-agent systems, there may be faulty or adversarial agents111 Note that our primary focus is on attacks at the agent layer, attacks can also occur at the communication layer [4, 5]. In the subsequent text, both faulty and adversarial agents are collectively referred to as faulty agents. against the rule to break the consensus of the system or drive the consensus to an unsafe value. To maintain the safe consensus of normal agents, the concept of resilient vector consensus has been developed. Resilient vector consensus means that all normal agents can achieve vector consensus where the final value lies within the convex hull of the initial states of normal agents despite the presence of faulty agents. Under resilient vector consensus, many distributed systems can work correctly under complex and changing circumstances. Therefore, the concept of resilient vector consensus is of paramount importance in the field of distributed systems.
In most consensus/resilient algorithms, agents directly send their states to their neighbors. However, if the transmitted states are intercepted by a malicious entity, this could lead to the disclosure of agents’ sensitive information, resulting in privacy leakage. A typical example is group decision-making [6]. For instance, in a voting process, a group of individuals aims to reach a unified opinion using consensus algorithms. Individuals can be reluctant to disclose their true opinions to others due to social pressure, fear of judgment, or other concerns [7]. Therefore, it becomes essential to protect the initial state of each participant, ensuring that their opinions remain confidential while still enabling the group to achieve consensus. This highlights the importance of privacy preservation as a critical aspect of resilient vector consensus.
@@
Due to the scarcity of communication and computation resources in distributed systems and the high demand of encryption algorithms on these resources, differential privacy has become the primary choice for privacy preservation in consensus algorithms. Typically, in resilient vector consensus, we aim to protect the initial state vectors of the normal agents without the existence of a trusted control center, where local differential privacy (local-DP) [8] is suitable to be utilized. Different from centralized differential privacy (central-DP), local-DP protects information during transmission from being intercepted, unrelated to differential attacks, and thus does not involve the concept of neighboring inputs. Local-DP also ensures that any pair of inputs generates similar outputs. However, the initial state range for each agent can be very large, and requiring pairs of distant initial states to produce similar outputs would reduce the system’s utility222 Utility refers to the extent to which a system can still perform its original functions after the application of privacy-preserving mechanisms. In this paper, we especially refer to the convergence accuracy.. This motivates us to use the concept of geo-privacy, characterizing the Euclidean distance between inputs as one of the metrics for privacy preservation. Specifically, we propose to use the concept of concentrated geo-privacy (CGP) [9], which is a generalization of differential privacy, to describe the privacy preservation of resilient vector consensus. Apart from a more precise description of privacy preservation, it also offers advantages such as the Gaussian mechanism and advanced composition.
@@
Much attention has been paid to resilient consensus algorithms, which can be roughly classified into two categories. One is to detect and preclude faulty agents[10, 11], while the other one is to try to find a safe value despite the presence of faulty ones[12, 13, 14, 15]. In this paper, we focus on the second one as it has more analytical results. Related works on resilient scalar/vector consensus and privacy-preserving resilient scalar consensus are provided below.
@@
Resilient scalar consensus is designed for normal agents to converge to a common state that lies between the maximum and minimum values of normal agents’ initial states. There are two frequently used ways: One way is to use the median to filter extreme values and reach resilience. In [16], Zhang et al. proposed a median consensus algorithm, where each normal agent updates by using its own value and the median of the states from its neighborhood. As for the other way, most of the algorithms rely on the strategy of simply disregarding suspicious values. For example, in a series of algorithms named Mean-Subsequence-Reduced (MSR) algorithm [12, 13], the main idea is to discard a fixed number of largest and smallest values in each normal agent’s neighborhood. Conversely, in the modification of the Weighted MSR (W-MSR) algorithm [14, 15], each normal agent only discards values that are greater or less than its own to maintain its state.
@@
Resilient vector consensus is ensured by that each normal agent must seek a point that is always located in the convex hull of its normal neighboring state vectors at each iteration. The existing work can be classified into three categories. The first category utilizes Tverberg partitions to compute Tverberg points. Specifically, it includes the Byzantine vector consensus algorithm [17] and Approximate Distributed Robust Convergence algorithm [2]. The second category calculates the intersection of multiple convex hulls. This involves Algorithm $1$ proposed in [18] and the multidimensional approximate agreement algorithm described in [19]. The algorithm of the last category utilizes the concept of “centerpoint”, exemplified by “Approximate Distributed Robust Convergence Using Centerpoint” (ADRC) algorithm [20]. However, these algorithms are rather time-consuming, and most of them require approximation algorithms, which consequently reduce fault tolerance. Notably, among these algorithms, the ADRC [20] stands out with its better tolerance for the mature approximation algorithm. In each iteration, every normal agent computes the centerpoint (an extension of the median in higher dimensions) of its neighborhood and updates its state as the weighted average of its current state and the centerpoint. Resilient vector consensus can be achieved as long as the communication topology and the number of faulty agents within the neighborhood meet certain requirements.
@@
Privacy-preserving resilient scalar consensus can primarily be achieved through three methods: cryptography-based methods [21], observability-based methods [22, 23], and noise-adding-based methods [21, 24]. Cryptography-based methods offer higher convergence accuracy; however, the processes of encryption, information exchange, and decryption are time-consuming, significantly reducing efficiency. Observability-based methods primarily reduce the observability of the system, making the system’s states unobservable to external entities. The most widely applied approach in this category is state decomposition. The remaining substates interact and update internally with the exchanged substate, thereby preserving privacy. However, this approach lacks complete privacy preservation and does not provide a quantifiable metric for the strength of privacy preservation. On the other hand, noise-adding-based methods are typically more flexible, as they allow for the adjustment of noise intensity and distribution based on specific application scenarios and privacy requirements. In [24], Fiore et al. proposed a Differentially Private MSR (DP-MSR) algorithm, which adds decaying and zero-mean Laplace noise to the scalar states of the normal agents. They also analyzed resilient scalar consensus in terms of differential privacy and asymptotics in expectation.
@@
Note that privacy-preserving resilient vector consensus is still an open issue. One simple way is to conduct a privacy-preserving resilient scalar consensus algorithm $d$ times, where $d$ is the dimension of states. However, this may cause the final value to fall outside the convex hull of the initial states of the normal agents [20]. Additionally, analyzing the performance of privacy-preserving resilient vector consensus by noise addition involves several key challenges: (i) The noise added makes the convergence nondeterministic, random, and asymptotic, rendering the existing definition of resilient vector consensus inapplicable. (ii) Due to the infinite sequence of noise and resilience, deriving an analytical solution for the final value is infeasible, making it difficult to quantify convergence accuracy. Additionally, high-dimensional convex hulls, unlike their one-dimensional counterparts defined by just two parameters, are characterized by numerous vertices and boundaries, which change randomly and sharply after noise is added. This randomness complicates both resilience and convergence accuracy analysis, particularly in arbitrarily high-dimensional settings. (iii) At each iteration, each agent transmits and updates its state information and the specific distribution of local state at each iteration is nondeterministic. We need to calculate the privacy leakage of the initial states caused by information transmission over infinitely many iterations, which is also a challenging task.
@@
To solve the above challenges, we consider a modification of ADRC, named Privacy-Preserving ADRC (PP-ADRC), and provide rigorous theoretical analysis for convergence and privacy performance. Compared to our conference version [25], we use Gaussian noise instead of Laplace noise and conduct a more rigorous performance analysis. Meanwhile, we add the final value distribution analysis by characterizing the deviation of the final value from the expectation. Moreover, we use CGP instead of DP to characterize the privacy preservation and compare it with DP comprehensively. We also add more simulations under a $3$-dimensional case to illustrate our theoretical results. The contributions of this paper are given as follows.
@@
  • We consider a modification of ADRC, named PP-ADRC, where each agent adds Gaussian noise to the local state vector for local interaction. We provide sufficient conditions to ensure resilient vector consensus in expectation.
@@
  • We employ $\rho$-CGP to characterize privacy preservation without detailed distributions of the outputs. Through a detailed comparison with differential privacy, we demonstrate that $\rho$-CGP can provide advantages for resilient vector consensus.
The organization of this work is as follows. We introduce preliminaries, problem formulation, and algorithm design in Sections 2 and 3, respectively. Then, we analyze the convergence condition and final value in Section 4. The privacy analysis is presented in Section 5. Simulation results are presented in Section 6. Finally, conclusions and avenues for future research are outlined in Section 7.
@@ 2 PRELIMINARIES
Basic Notations: Throughout this paper, we denote by $\mathbb{R}^{d}$ the $d$-dimensional Euclidean space, by $\mathbb{R}^{m\times d}$ the space of all $m\times d$-dimensional matrices, and by $(\mathbb{R}^{m\times d})^{\mathbb{N}}$ the space of matrix-valued sequences in $\mathbb{R}^{m\times d}$, where $\mathbb{N}$ denotes the set of natural numbers. The natural logarithm is denoted by $\log$. Let $\mathbf{A}$ be a matrix. Then, we denote by $[\mathbf{A}]_{i},[\mathbf{A}]^{j},[\mathbf{A}]_{ij},\mathbf{A}^{\top}$ the $i$ th row, the $j$ th column, the $(i,j)$ th element, and the transpose of the matrix $\mathbf{A}$. We then consider two matrices $\mathbf{A}$ and $\mathbf{B}$, and if $[\mathbf{A}]_{ij}\leq[\mathbf{B}]_{ij}$ holds for any pair of $i,j$, we say $\mathbf{A}\leq\mathbf{B}$. For a vector $x\in\mathbb{R}^{d}$, we denote by $x_{\operatorname{max}}$ and $x_{\operatorname{min}}$ maximum element and minimum element of it, respectively. Then, we consider a matrix consisting of $n$ row vectors $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]^{\top}$, where $x_{i}$ denotes the $i$th vector and $x_{i,k}$ stands for the $k$th element of the vector $x_{i}$. We denote by $\mathbf{1}$ a column vector of ones, and by $I_{d}$ the $d$-dimensional identity matrix. A matrix is row-stochastic if all its elements are nonnegative and each row sums up to $1$. For a given point set $C\subseteq\mathbb{R}^{d}$, we denote by $\text{conv}(C)$ the convex hull of the point set $C$, by $|C|$ the cardinality of $C$. For two different vectors $x,x^{\prime}\in\mathbb{R}^{d}$, we represent by $\mathrm{dist}(x,x^{\prime})=||x-x^{\prime}||_{2}$ the Euclidean distance between $x$ and $x^{\prime}$. For two $n$-tuples of vectors $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]^{\top}$ and $\mathbf{x}^{\prime}=[x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{n}^{\prime}]^{\top}$, $\mathrm{dist}(\mathbf{x},\mathbf{x}^{\prime})$ stands for the maximum distance among all the vectors, which is $\max\limits_{i}\operatorname{dist}(x_{i},x_{i}^{\prime})$. In probability theory, given a random variable $z\in\mathbb{R}$, we denote by $\mathbb{E}(z)$ and $\operatorname{var}(z)$ the expectation and the variance of $z$, respectively. For a random vector $z=[z_{1},\ldots,z_{d}]^{\top}\in\mathbb{R}^{d}$, the expectation of $z$ is $\mathbb{E}(z)=[\mathbb{E}(z_{1}),\ldots,\mathbb{E}(z_{d})]^{\top}$. For a topological space $\mathcal{T}$, we denote by $\mathcal{B}(\mathcal{T})$ the set of Borel subsets of $\mathcal{T}$, and $\mathrm{P}$ is the corresponding probability measure. For a zero-mean $d$-dimensional Gaussian distribution, the probability density function (pdf) is given by $G(x;\Sigma)=\frac{1}{\sqrt{(2\pi)^{d}|\Sigma|}}\mathrm{exp}\left(-\frac{1}{2}x^{\top}\Sigma^{-1}x\right)$. For two functions $f(x)$ and $g(x)$, $f(x)=\Omega(g(x))$ means that $g(x)$ is the asymptotic lower bound of $f(x)$.
@@
Privacy Notions: We first give two formal definitions of the standard local-DP and then introduce CGP.
@@
($\varepsilon$-DP): Given spaces $U$, $V$, and $\varepsilon\in\mathbb{R}\geq 0$, a randomized function $M:U\rightarrow V$ is $\varepsilon$-differentially private, if for any pair of inputs $x$, $x^{\prime}\in U$ and any $S\subseteq V$, we have $\mathrm{P}\{M(x)\in S\}\leq e^{\varepsilon}\mathrm{P}\{M(x^{\prime})\in S\}.$
@@
($(\varepsilon,\delta)$-DP): Given spaces $U$, $V$, and $\varepsilon$, $\delta\in\mathbb{R}\geq 0$, a randomized function $M$ is $(\varepsilon,\delta)$-differentially private, if for any pair of inputs $x$, $x^{\prime}\in U$ and any $S\subseteq V$, $\mathrm{P}\{M(x)\in S\}\leq e^{\varepsilon}\mathrm{P}\{M(x^{\prime})\in S\}+\delta$ holds.
DP introduces randomness into the output, making it nondeterministic and following a probability distribution. $\varepsilon$-DP ensures that the probability ratio of obtaining the same result for any two inputs is at most $e^{\varepsilon}$. In contrast, $(\varepsilon,\delta)$-DP relaxes this, allowing the ratio to exceed $e^{\varepsilon}$ with a probability no greater than $\delta$, i.e., $\varepsilon$-DP is achieved with a probability of at least $1-\delta$. Note that in both definitions, the inputs $x$ and $x^{\prime}$ can be in the real domain and consist of one or more scalars or vectors. The parameter $\varepsilon$ in both definitions characterizes the privacy preservation strength, with smaller values of $\varepsilon$ providing stronger privacy preservation.
Additionally, the Gaussian mechanism is only applicable to $(\varepsilon,\delta)$-DP, due to the light tail of the Gaussian distribution. The pdf ratio for two Gaussian distributions with different means tends to infinity, making it impossible to bound the ratio by $e^{\varepsilon}$. In contrast, the Laplace mechanism, with a bounded probability density ratio for different means, is suitable for $\varepsilon$-DP. For further details, we refer to [26].
It is worth noting that Definitions $3$-$4$ impose rather strict requirements for resilient vector consensus. As the range of possible initial states of an agent is typically large, ensuring that two far apart initial states produce similar outputs would cause significant system utility reduction. To solve this issue, geo-privacy has been widely studied, which uses Euclidean distance as one metric for privacy preservation. Supporting the Gaussian mechanism and better composition, CGP is selected here, which is established on the Rényi divergence given below [27].
@@
Note that $D_{\alpha}(F\|G)$ is determined by the expectation of the $\alpha-1$ power of the ratio of two distributions at each point, quantifying the difference between two distributions. Meanwhile, larger values of $\alpha$ emphasize parts where the ratio is greater, while smaller values of $\alpha$ focus more on the overall average differences. The roles of logarithm and $\frac{1}{\alpha-1}$ are intended to prevent numerical overflow in the results. Rényi divergence is monotonically non-decreasing with $\alpha$.
@@
($\rho$-CGP): Given spaces $U$, $V$, and $\rho\in\mathbb{R}\geq 0$, a randomized function $M:U\rightarrow V$ is said to satisfy $\rho$-CGP, if for any inputs $x,x^{\prime}\in U$ and all $\alpha>1$, it holds that
@@
In CGP, $\rho$ provides the upper bound on $D_{\alpha}(M(x)\|M(x^{\prime}))/\alpha/\mathrm{dist}(x,x^{\prime})^{2}$, characterizing the difference between the output distributions of $M$ for any pair of inputs, thereby quantifying the strength of privacy preservation. A smaller $\rho$ indicates that $M$ is less sensitive to the inputs, thereby providing stronger privacy preservation. Besides, the privacy should degrade gracefully as $\mathrm{dist}(x,x^{\prime})$ increases. The parameter $\alpha$ plays a crucial role as it allows for the adjustment of privacy preservation strictness. A larger $\alpha$ emphasizes the worst-case scenario by highlighting the regions where the two distributions differ the most, while a smaller $\alpha$ focuses on the average difference. This flexibility enables the adaptation of privacy preservation levels according to varying scenario requirements.
Under local-DP, the privacy parameter $\varepsilon$ provides a uniform upper bound based on the maximum distance between inputs, often resulting in stronger privacy for most input pairs than $\varepsilon$ indicates. In contrast, CGP integrates input distance into its privacy analysis, offering tighter, distance-dependent privacy guarantees. This allows CGP to achieve equivalent privacy preservation with less noise, thereby improving system utility.
3 Problem Formulation and Algorithm Design
In this section, we first present the network model of the multi-agent system with faulty agents and define privacy attackers. We then introduce the PP-ADRC algorithm and provide a brief explanation of centerpoint and its computation methods. Finally, we state several interesting problems.